### v1

# Overcoming Data Scarcity in Generative Language Modelling for Low-Resource Languages: A Systematic Review

Abstract.
Generative language modelling has surged in popularity with the emergence of services such as ChatGPT and Google Gemini. While these models have demonstrated transformative potential in productivity and communication, they overwhelmingly cater to high-resource languages like English. This has amplified concerns over linguistic inequality in natural language processing (NLP). This paper presents the first systematic review focused specifically on strategies to address data scarcity in generative language modelling for low-resource languages (LRL). Drawing from 54 studies, we identify, categorise and evaluate technical approaches, including monolingual data augmentation, back-translation, multilingual training, and prompt engineering, across generative tasks. We also analyse trends in architecture choices, language family representation, and evaluation methods. Our findings highlight a strong reliance on transformer-based models, a concentration on a small subset of LRLs, and a lack of consistent evaluation across studies. We conclude with recommendations for extending these methods to a wider range of LRLs and outline open challenges in building equitable generative language systems. Ultimately, this review aims to support researchers and developers in building inclusive AI tools for underrepresented languages, a necessary step toward empowering LRL speakers and the preservation of linguistic diversity in a world increasingly shaped by large-scale language technologies.
Generative Language Modelling, Low-resource Languages, Data Scarcity, Systematic Review, Data Augmentation, PRISMA
$c$s: Computing methodologies Natural language generationccs: Computing methodologies Language resources
1. Introduction
Language modelling is a central task in the field of natural language processing (NLP) that aims to capture the features of a language by extracting the statistical and contextual relationships within a language (Wang et al., 2023). This information can subsequently be used to inform next-word predictions for a given sequence of previous terms (Kłosowski, 2018). It involves enlisting a model to extract the characteristics of a language from a large corpus of textual data that can consequently be used to process, understand and generate text (Thirunavukarasu et al., 2023). Language models evolved from statistical methods like Markov processes (Almutiri and Nadeem, 2022) and n-gram models (Roark et al., 2007) to recurrent neural network (RNN) approaches, such as Long Short-Term Memory (LSTM) RNNs (Van Houdt et al., 2020). However, the current rise in popularity of language models and language model-powered services such as ChatGPT (Lian et al., 2024; Ng and Chow, 2024) can be attributed to the breakthrough that introduced the transformer (Vaswani, 2017) as an architecture for modelling languages. This advancement enabled the development of pre-trained models such as the transformative Generative Pretrained Transformer (GPT) (Radford, 2018) and Bidirectional Encoder Representations from Transformers (BERT) (Devlin, 2018), with the transformer architecture enhancing NLP tasks such as machine translation (Raganato and Tiedemann, 2018), classification (McGiff and Nikolov, 2024) and named entity recognition (Arkhipov et al., 2019).
Generative language modelling is a subset of language modelling that focuses on building models that can produce new sentences for a set of given semantics (Su et al., 2019). Generative language modelling is associated with the topic of natural language generation (NLG) as natural language text can be produced for a variety of communicative goals (Dong et al., 2022). This differs from natural language understanding (NLU) whereby the morphology, syntax and semantics are processed to comprehend the meaning of text, identifying intent or extracting information (Xu et al., 2020). Generative language models are the driving force behind popular services such as ChatGPT (Lian et al., 2024; Ng and Chow, 2024), as they process a prompt and produce natural language text that is aligned with the context of the input. These models are flexible in their ability to represent several languages at once, with studies successfully modelling both individual languages (De Mattei et al., 2020; Müller and Laurent, 2022; Simoulin and Crabbé, 2021) and groups of languages (Liu et al., 2020; Xue et al., 2021).
NLP research has been dominated by a minority of the world’s languages (Nigatu et al., 2024; Joshi et al., 2020; Bird, 2022). Nigatu et al. describe the ’low-resource’ and ’high-resource’ categories that the NLP community has adopted to classify the research interest for a given language (Nigatu et al., 2024). Their study echoes the reality that low-resource languages (LRL) are constantly trailing the NLP developments and breakthroughs experienced by high-resource languages. Although ”low-resource” can be interpreted as less studied and less computerised (Magueresse et al., 2020), it can also be used to describe a language with scarce resources (Ogueji et al., 2021). Some evidence suggests that the availability of NLP researchers with fluency in a language correlates with the availability of datasets (Yu et al., 2022). This connects the theoretical challenge of sourcing data for language modelling with the challenge of ensuring that LRL speakers are not left behind in technological advancements. Although these LLM-based services can improve worldwide communication and information access, their focus on building and optimising their applications for English could hinder linguistic diversity (Zhang et al., 2023; Kirk et al., 2023). In terms of language preservation, some research has connected generative language models with being capable of revitalising endangered languages (Nanduri and Bonsignore, 2023) and conserving languages that are less predominantly modelled by the major LLM providers (Hutson et al., 2024; Lai et al., 2023). Therefore, developing a generative language model that accurately captures the linguistic features of an LRL could serve as a digital preservation tool, mitigating the risk of its extinction. The need to support LRLs with modern technological advancements has motivated some research groups that are working towards improving LRL representation in NLP and NLG research (Singh et al., 2024; Team et al., 2024; Le Scao et al., 2023).
A large volume of data is required to build a robust generative language model (Kaplan et al., 2020; Muennighoff et al., 2023; Hoffmann et al., [n. d.]; Radford et al., 2019). This conflicts with the limited textual data that is available to LRLs. Many studies aim to tackle the data disparity, with one main branch of these approaches focusing on integrating a data augmentation process into the model training pipeline (Shorten et al., 2021). Data augmentation is a process that generates synthetic data from an existing dataset (Shorten and Khoshgoftaar, 2019). The concept of data augmentation is commonly used in the field of computer vision to train more robust models (Taylor and Nitschke, 2018; Mikołajczyk and Grochowski, 2018; Khosla and Saini, 2020). Image datasets are often enlarged by performing image transformations such as rotations or brightness modifications to create additional training images (Mumuni and Mumuni, 2022). While it has been argued that NLP has not adopted data augmentation to the same degree as computer vision, a wide range of studies have enlisted a variety of processes to bridge the gap in available training data for LRL modelling (Shorten et al., 2021). Although prior surveys have explored data augmentation in NLP broadly, no systematic review to date has focused on how these strategies are applied in the development of generative models for LRLs.
To address this gap, this paper presents the first systematic review focused specifically on techniques for overcoming data scarcity in the context of generative language modelling for low-resource languages. By analysing 54 studies, this review provides a structured overview of the technical approaches used, the language families studied, and the model architectures employed. It also highlights underrepresented languages, evaluates the effectiveness of key methods, and identifies research gaps in the development of generative models for LRLs. Therefore, the objectives of this systematic review are to:
  • Identify languages and language families in particular that are underrepresented in this area of LRL modelling.
  • Determine the methods that have been used to mitigate data scarcity, including but not limited to data augmentation in the development of generative language models for LRLs.
  • Analyse the impact of methods for overcoming data scarcity on the performance of various language modelling tasks.
  • Identify existing gaps in the literature and share insights relating to future research directions in generative LRL modelling.
2. Review Methodology
A systematic review is described by Page et al. as being a review that enlists clear, methodical techniques to aggregate and summarise the results of studies aimed at answering a clearly defined research question (Page et al., 2021). It is an informative process that can yield in-depth exploration and determine the current state of existing knowledge in a field. The PRISMA 2020 statement indicates that a systematic review can combine information relating to future research opportunities, answer research questions that span different individual studies, identify shortcomings in research and explore existing primary research results (Page et al., 2021). The PRISMA review methodology can be divided into four steps for this review: (1) Defining the research objectives/questions and eligibility criteria; (2) Adopting a search strategy; (3) Performing the study selection; and (4) Data extraction.
2.1. Defining the Research Questions and Eligibility Criteria
This systematic review focuses on the main research objective of creating generative language models for LRLs using methods to mitigate the challenges posed by limited data. However, we have identified several research sub-questions that will help to inform this overarching research objective.
Defining explicit criteria for including or excluding evidence is necessary for enabling readers to both understand the scope of the review and to verify eligibility decisions (Page et al., 2021). We apply the Population, Intervention, Comparison, and Outcome (PICO) framework to our research objective to define the eligibility criteria for evidence in this systematic review (Stone, 2002). The PICO framework consists of four key elements: Population (P), which defines the target group under study; Intervention (I), which refers to the treatment, technique, or approach being evaluated; Comparison (C), which contrasts the intervention against an alternative or baseline; and Outcome (O), which represents the measurable effects or improvements resulting from the intervention. The PICO framework is employed to formulate a well-defined research question, which serves as the foundation for constructing a systematic database search string in accordance with the PRISMA methodology. LRLs are considered to be the population of the framework. On the other hand, the techniques for overcoming data scarcity are considered to be the intervention element of the framework. The comparison section of the framework is linked to comparing the performance of generative language models, both when using data augmentation or other technical methods to address data disparity for LRLs, and when not employing these methods. The outcome element of the framework relates to quantitative (eg: sacreBLEU, COMET etc) and qualitative (eg: human feedback) improvement in language generation. The PICO framework produces the following revised research question: In the context of LRLs (Population), how do techniques designed to overcome data scarcity (Intervention) compare to traditional NLG methods (Comparison) in improving the performance of generative language models (Outcome)?
Once the revised research question was formulated using the PICO framework, we devised a set of inclusion and exclusion criteria to filter out the most irrelevant studies from our search. We included the papers that satisfied at least one of the following inclusion criteria:
  1. IC1.
    The paper focuses on building generative language models for LRLs.
  2. IC2.
    The paper enlists a technical method for overcoming data scarcity.
  3. IC3.
    The paper reports their methods or reviews existing methods that tackle generative language modelling for LRLs.
Additionally, we excluded the papers that satisfied at least one of the following exclusion criteria:
  1. EC1.
    Papers not written in English.
  2. EC2.
    We excluded all papers not published in journals, conference proceedings, workshops, symposia, or as arXiv preprints, and we also excluded posters and abstracts.
  3. EC3.
    Papers that create language models for NLP-specific tasks as opposed to NLG tasks.
  4. EC4.
    Papers published before 2015, or preprints released before 2021.
  5. EC5.
    Papers that do not produce results or survey existing literature relating to creating NLG models for LRLs or data augmentation techniques.
2.2. Adopting a Search Strategy
In terms of the search strategy for this systematic review, we initially compiled a list of information sources for the search itself. The strategy involved searching SpringerLink, ACL Anthology, ACM, IEEE Xplore, Scoupus, Wiley Interscience and Google Scholar. These platforms were chosen for their expansive digital repositories. Specific sources, such as the ACL Anthology, were included due to their relevance to the computational linguistics focus of the research objective. Table 1 details the search strings that were used for the various digital repositories. It is worth flagging that some of the platforms differ in terms of the syntax required for building search strings. Therefore, we report the equivalent strings for each platform. The searches match elements in the search string with metadata associated with each study such as the title, keywords and abstract. Most search engines use a combination of logical OR and logical AND operations to combine different terms and to impose constraints on groups of terms respectively. Search string keywords were chosen to match the main terms in the PICO research objective and sub-questions. Synonyms for terms such as ”low resource language” and ”language model” were sourced and included in the search string to further ensure the retrieval of relevant studies.
Table 1. Search string composition as part of the search strategy.
Scope String
Generative Aspect ”generative”
Model Type ”language model” OR ”text model”
Language Type ”low resource language” OR ”minority language” OR ”endangered language”
Exclusion ”classification”
2.3. Performing the Study Selection
The study selection phase of this research was initially carried out using ASReview, an open-source machine learning-powered tool for performing efficient systematic reviews (Van De Schoot et al., 2021). The study was subsequently completed with a manual review of the remaining papers. ASReview assists researchers by predicting the most relevant studies and ranking them accordingly. We found that this accelerated the study selection process. The ASReview element of the selection process involved loading a dataset of bibliographic records from the previously listed information sources in RIS format. Once the dataset of papers was successfully imported, we chose TF-IDF for feature extraction, Naive Bayes as the classifier, ASReview’s maximum query strategy and dynamic resampling as the balance strategy. The systematic review tool begins the screening process by displaying a random sample of studies to initialise the active learning element. Studies can be labelled as being relevant or irrelevant with these classifications being used to predict the next most relevant paper. The papers that are most likely to be relevant are prioritised for review. It is necessary to report that a single reviewer worked independently on each stage of the screening process. This process was used for the first pass through the original dataset of papers collected from the previously defined information sources. The first pass removed all papers that satisfied EC1, EC3 and EC4. The second pass involved screening papers that satisfied EC2 or EC5. This involved manually reading the remaining papers and adding relevant papers found via reference tracking to the list of included studies.
2.4. Data Extraction
The data collection process involved extracting information relating to qualitative and quantitative features of the studies that align with the objective and research questions of this systematic review. The attributes extracted from the studies are formulated into data items represented via data rows in a spreadsheet. In terms of the quantitative information, the data items included: the title of the paper; whether it is a journal, conference, workshop, symposium or preprint paper; the publication name and a list of results for quality assessment questions. The quality assessment questions were formulated following PRISMA guidelines and existing review papers (Peixoto et al., 2021; Zapata et al., 2015; Sardi et al., 2017). The purpose of the quality assessment questions is to help gauge the relevance and value of the studies in relation to the research objective of the systematic review. We report the following quality assessment questions:
  1. QA1.
    Does the paper discuss building language models for LRLs? (+1) Yes/ (+0) No
  2. QA2.
    Does the paper present empirical results? (+1) Yes/ (+0) No
  3. QA3.
    Does the paper discuss a process for overcoming data scarcity? (+1) Yes/ (+0) No
  4. QA4.
    Has the paper presented and compared the results of using data augmentation? (+1) Yes/ (+0) No
  5. QA5.
    Was the paper published in a relevant journal or conference proceedings? Given that there are different metrics for categorising the relevance of journal and conference publications, we consulted the Journal Citation Reports 2024 and CORE 2023 to determine a suitable score for each study. For conferences, studies were scored based on the CORE 2023 ranking: +1.5 for CORE A, +1 for CORE B, +0.5 for CORE C, and +0 for unranked conferences. For journals, scores were assigned using the Journal Citation Reports 2024: +2 for Q1, +1.5 for Q2, +1 for Q3/Q4, and +0 for unranked journals. Other publications received a score of +0.
In terms of the qualitative information, the data items included attributes representing each of the research sub-questions that were formulated to better inform the overall research objective of this systematic review. We report the following qualitative research sub-questions for data extraction:
  1. RQ1:
    What low-resource languages have already been modelled by other studies? This should consist of a list of languages that the study models. While some studies could model a single language, others could model a range of languages.
  2. RQ2:
    What technical methods have been used to overcome data scarcity in building generative language models? Answers for this question should represent categories relating to specific technical methods.
  3. RQ3:
    What text data augmentation techniques have been used to build generative language models for low-resource languages? This section should include all the text data augmentation techniques that produce new data from existing textual data.
  4. RQ4:
    Which sources are publishing in the area of low-resource language modelling? This attribute should consist of the name of the publisher of the study.
  5. RQ5:
    What architectures are being used to produce low-resource language models? This research question should produce a list of the model architectures that are used by the study to model LRLs.
  6. RQ6:
    How effective have methods for overcoming data scarcity been at improving model performance for different NLG tasks? This attribute should include various metrics that were produced to evaluate the NLG models.
  7. RQ7:
    How scarce is the data in studies tackling low-resource language modelling? This should consist of a quantity that describes the size of the dataset being used for training generative language models for LRLs.
  8. RQ8:
    How are the models in these studies being evaluated? This research question should explore the evaluation methods that are being used to measure the performance of generative language models for LRLs.
  9. RQ9:
    What is the NLG task? This section should include the categories of NLG tasks that the study tackles for LRLs.
3. Results
This section is structured into two parts: the first presents the outcomes of the study selection process and the second details the findings related to the extraction of research question-specific characteristics from the included studies.
3.1. Study Selection
The study selection section of this systematic review was split into three phases. A set of 642 papers were found by searching the previously mentioned information sources. Performing deduplication on the dataset removed 52 papers from the set and left 590 studies for screening. The first pass using ASReview for the screening process resulted in the rejection of 486 papers based on their title, keywords, abstract and whether they satisfy EC1, EC3, or EC4. The first pass retained 104 of the studies. The second pass screened for papers that satisfied EC2 or EC5. This second pass rejected 55 papers, retaining 49 papers in total. However, five papers were identified via reference tracking and consequently brought the final total to 54 papers for review. Figure 1 represents the PRISMA flow diagram that reflects the screening process used to select the studies to include in this systematic review.
The 54 selected studies were subsequently evaluated using the predefined quality assessment criteria. The results of applying quality assessment are illustrated in Figures 24. Table 2 reports the quality assessment score assigned to each paper. The majority of papers in this review (77.78%) tackle building language models for a LRL. 88.89% of the literature present empirical results. Additionally, 75.93% of studies discuss the use of a process for overcoming data scarcity. Over half of the studies (51.85%) present and compare methods aimed at addressing data scarcity. Among the conference papers, 28.6% were published in A* venues, 14.3% in A-ranked and B-ranked venues respectively, 9.5% in C-ranked venues, and 33.3% were unranked, based on CORE conference rankings. Among the journal papers, 15.8% were classified as Q1, 21.1% as Q2, 57.9% as Q3, and 5.3% as Q4, based on the 2024 Journal Impact Factor quartile rankings.
Figure 1. This PRISMA flow diagram describes the process of using the PICO framework-based search string to find and filter relevant studies. After applying each exclusion criteria and including additional papers via reference tracking, 54 papers were included in the systematic review.
Figure 2. Distribution of papers based on whether they tackle building a language model for a low-resource language (QA1), and whether they present empirical results (QA2).
Figure 3. Distribution of papers based on whether they discuss methods for addressing data scarcity (QA3), and whether they present and compare results of such methods (QA4).
Figure 4. CORE ranking distribution for conference papers and Journal Impact Factor 2024 distribution for journal papers (QA5).
Table 2. Overview of data extracted for each quality assessment criterion: QA1–QA5. QA1: Does the paper discuss building language models for low-resource languages (LRLs)? QA2: Present empirical results? QA3: Discuss a method for overcoming data scarcity? QA4: Present and compare data augmentation results? QA1–QA4 are each scored as +1 for Yes, +0 for No. QA5 assesses publication venue quality: conferences are scored using the CORE 2023 ranking (+1.5 for CORE A, +1 for B, +0.5 for C, +0 for unranked); journals follow Journal Citation Reports (JCR) 2024 (+2 for Q1, +1.5 for Q2, +1 for Q3/Q4, +0 for unranked).
Paper Publication Year Publication Type QA1 QA2 QA3 QA4 QA5 Total Score
(Li et al., 2024b) 2024 conference y y y y A 5.50
(Sun et al., 2019) 2019 journal y y y y Q2 5.50
(Zhang et al., 2020) 2020 conference y y y y A* 5.50
(Usui and Komiya, 2023) 2023 conference y y y y A 5.50
(Sorokin et al., 2022) 2022 conference y y y y A 5.50
(Gao et al., 2019) 2019 conference y y y y A* 5.50
(Bhowmick et al., 2023) 2023 journal y y y y Q3 5.00
(Otegi et al., 2020) 2020 conference y y y y B 5.00
(Marie and Fujita, 2020) 2020 journal n y y y Q1 5.00
(Ahmadnia and Dorr, 2019) 2019 journal y y y y Q3 5.00
(Yirmibeşoğlu and Güngör, 2023) 2023 journal y y y y Q3 5.00
(Maimaiti et al., 2021) 2021 journal y y y y Q3 5.00
(Cahyawijaya et al., 2021) 2021 conference y y y n A* 4.50
(Guo et al., 2021a) 2021 conference y y y y A* 4.50
(Wongso et al., 2023) 2023 journal y y y n Q2 4.50
(Dandapat and Federmann, 2018) 2018 conference y y y y C 4.50
(Wang et al., 2018) 2018 conference n y y y A* 4.50
(Gao et al., 2020) 2020 conference n y y y A* 4.50
(Pham et al., 2021) 2021 preprint y y y y - 4.00
(Niyogi and Bhattacharya, 2024) 2024 preprint y y y y - 4.00
(Downey et al., 2024) 2024 preprint y y y y - 4.00
(Hong et al., 2024) 2024 preprint y y y y - 4.00
(Chang et al., 2023) 2023 preprint y y y y - 4.00
(Shafayat et al., 2024) 2024 preprint y y y y - 4.00
(Tanwar and Majumder, 2020) 2020 journal y y y n Q3 4.00
(Gerz et al., 2018) 2018 journal y y n n Q1 4.00
(Berckmann and Hiziroglu, 2020) 2020 conference y y y y - 4.00
(Liu, 2022) 2022 workshop y y y y - 4.00
(Pham Van and Le Thanh, 2022) 2022 symposium y y y y - 4.00
(Nissanka et al., 2020) 2020 conference y y y y - 4.00
(Agarwal et al., 2022) 2022 workshop y y y y - 4.00
(Scalvini and Debess, 2024) 2024 conference y y y n B 4.00
(Guo et al., 2024b) 2024 conference y y y n B 4.00
(Ulčar and Robnik-Šikonja, 2023) 2023 journal y y n n Q2 3.50
(Bendel and Jabou, 2024) 2024 journal y y n n Q2 3.50
(Wu and Wang, 2023) 2023 conference n y y y - 3.00
(Babaali et al., 2024) 2024 journal y y n n Q4 3.00
(Li et al., 2022a) 2022 journal y y n n Q3 3.00
(Wang et al., 2022) 2022 journal n n y n Q1 3.00
(Mi et al., 2024) 2024 journal y y n n Q3 3.00
(Yu et al., 2020) 2020 journal n y y n Q3 3.00
(Holiel et al., 2023) 2023 conference n y y n C 2.50
(Mao and Yu, 2024) 2024 workshop y y n n - 2.00
(Toraman, 2024) 2024 preprint y y n n - 2.00
(Zhang et al., 2024a) 2024 preprint y y n n - 2.00
(Acikgoz et al., 2024) 2024 preprint y y n n - 2.00
(Haque et al., 2021) 2021 journal n n y n Q3 2.00
(Shi et al., 2022) 2022 journal n n y n Q3 2.00
(Zhang et al., 2024b) 2024 journal n n y n Q3 2.00
(Salim et al., 2023) 2023 conference y y n n - 2.00
(Saini and Vidhyarthi, 2023) 2023 conference y y n n - 2.00
(Rizqullah et al., 2023) 2023 conference n y y n - 2.00
(Agarwal et al., 2023) 2023 conference y n n n - 1.00
(Wang et al., 2021) 2021 preprint n n y n - 1.00
3.2. RQ1: What Low-Resource Languages Have Already Been Modelled by Other Studies?
This review has identified a wide range of LRLs that have been modelled by other studies. A total of 295 unique LRLs have been trained spanning a variety of language generation tasks and different model architectures. Figure 5 showcases the most frequently modelled LRLs by studies examined in this review. Bengali is the most frequently modelled LRL in this review, with 13% of papers including it (Niyogi and Bhattacharya, 2024; Salim et al., 2023; Chang et al., 2023; Shafayat et al., 2024; Bhowmick et al., 2023; Sorokin et al., 2022; Agarwal et al., 2022). Hindi follows behind Bengali with the language appearing in 11% of the studies (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Bhowmick et al., 2023; Tanwar and Majumder, 2020; Gerz et al., 2018; Maimaiti et al., 2021). Whereas, Tamil (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Gerz et al., 2018; Nissanka et al., 2020), Telugu (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Sorokin et al., 2022; Tanwar and Majumder, 2020; Dandapat and Federmann, 2018) and Turkish (Chang et al., 2023; Agarwal et al., 2022; Gerz et al., 2018; Toraman, 2024; Acikgoz et al., 2024; Yirmibeşoğlu and Güngör, 2023) are each modelled by 9% of the papers, thus making them the third most commonly modelled LRLs. Although Turkish has been considered to be a high-resource language (Chang et al., 2023), some papers argue it is an LRL (Toraman, 2024; Yirmibeşoğlu and Güngör, 2023). Therefore we classify Turkish as an LRL in this review. Moreover, Javanese (Chang et al., 2023; Gerz et al., 2018; Cahyawijaya et al., 2021; Wongso et al., 2023) and Khmer (Agarwal et al., 2022; Gerz et al., 2018; Mao and Yu, 2024; Pham Van and Le Thanh, 2022) have both appeared in 7% of the papers, with Azerbaijani (Chang et al., 2023; Pham et al., 2021; Maimaiti et al., 2021), Belarusian (Chang et al., 2023; Mao and Yu, 2024; Pham et al., 2021) and Slovak (Pham et al., 2021; Chang et al., 2023; Gerz et al., 2018) among the 16 languages that were modelled in 6% of studies times. Figure 5 is a word cloud that displays more of the most frequently modelled LRLs in this review. This word cloud highlights that LRLs such as Marathi (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Saini and Vidhyarthi, 2023) and Odia (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Agarwal et al., 2023) were also modelled somewhat frequently. Figure 6 reveals aggregations whereby each language is assigned to a category according to its respective language family. We have identified languages from 39 different language families that have been modelled by studies in this review. LRLs belonging to the Indo-European language family consist of 35% of the most frequently modelled LRLs. Austronesian languages are the second most commonly modelled LRL with a frequency of 15% in this review. Turkic and Niger-Congo-based languages are similarly represented with 8% and 7% frequency respectively. Uralic, Afro-Asiatic and Sino-Tibetan each represent 4% of the LRL families modelled by the literature in this review. Six remaining language families are represented with a frequency greater than 1%. These language families include Creole, Atlantic-Congo, Austro-Asiatic and Nakh-Daghestanian at 2%, with Mongolic and Constructed at 1%.
Figure 5. Visualisation of the most frequently modelled low-resource languages (RQ1), shown as a word cloud and a bar chart of the top 10 LRLs.
Figure 6. Distribution of the most frequently modelled low-resource language families by papers in this systematic review (RQ1).
3.3. RQ2: What Technical Methods Have Been Used to Overcome Data Scarcity in Building Generative Language Models?
A wide variety of technical methods have been used to overcome the issue of data scarcity associated with building generative language models for LRLs. We created a variety of categories that reflect the range of technical approaches undertaken by these papers that relate to filling the data gap required to create generative models. These categories are not mutually exclusive, as many papers employed multiple strategies in parallel. Each paper was assigned to one or more categories based on the techniques it discussed or implemented to tackle the issue of data scarcity. Figure 7 indicates that the most frequently mentioned technical approach was to perform some form of augmentation on monolingual data. The majority of the papers reference augmenting monolingual data, back-translation, building multilingual models, building monolingual models or prompting existing generative language models to approach the challenge of limited data for training. Table 3 provides an overview of how individual papers align with specific technical methods for overcoming data scarcity. The following sections discuss these methods in order of frequency, starting with the most commonly used approaches.
3.3.1. Monolingual Text Data Augmentation
The application of various strategies for bolstering monolingual text data occurs in 26% of the studies in this review. Such methods include generating new sentences for neural machine translation (NMT) by substituting low-frequency words with terms selected by an LSTM language model (Haque et al., 2021; Shi et al., 2022); constructing new sentences through the grammatical reordering of monolingual data using a reverse transcription grammar model (Wu and Wang, 2023); using a generative model combined with a grammar correction model to create grammatically correct synthetic data from the monolingual source data (Sun et al., 2019); and duplicating original sentences in a monolingual corpus and using a model to paraphrase them (Shi et al., 2022; Guo et al., 2021a; Gao et al., 2020). Other approaches include synthesising user-generated text for NMT by applying noise and stylistic changes to textual data (Marie and Fujita, 2020); appending document-specific metadata and information to the document itself (Usui and Komiya, 2023); generating text data for NMT by turning the process of augmentation into an optimisation problem (SwitchOut) (Haque et al., 2021; Shi et al., 2022; Wang et al., 2018); and using POS tagging to make replacements based on the label of the term (Maimaiti et al., 2021; Shi et al., 2022). Additional strategies involve boosting training data by reordering monolingual target language text to align with the structure of another language (Haque et al., 2021); incorporating knowledge at the word, phrase, and sentence levels to maximise the information gain from a monolingual corpus (Mi et al., 2024); and using soft contextual data augmentation to randomly replace a word with a similar meaning (Haque et al., 2021; Gao et al., 2019).
3.3.2. Back-Translation
Back-translation is also frequently used as a technique for addressing data scarcity, appearing in 24% of the studies in this review (Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Pham et al., 2021; Shi et al., 2022; Marie and Fujita, 2020; Wang et al., 2018; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Wang et al., 2022). This approach involves boosting monolingual data by using a backward NMT model to translate a monolingual corpus from a target language to a source language. This process of translating to another language and back again can produce variations that bolster the source language dataset. Although this technique starts with monolingual data and shares some characteristics with monolingual augmentation methods, this review treats back-translation as a distinct category due to its reliance on machine translation systems.
3.3.3. Prompt Engineering
A considerable portion of the studies in this review, specifically 20%, focused on addressing data scarcity by prompting existing generative language models. These existing models are examined based on their capacity to generate text for a wide range of LRLs. Several techniques have been proposed to improve generative language models. The programmer-interpreter approach iteratively refines a model’s output using a smaller language model for modifications (Li et al., 2024b). Another approach involves few-shot learning, where a model is taught unseen languages through examples and definitions (Rizqullah et al., 2023). Retrieval-augmented generation, which incorporates textbook-like material, enhances prompts and improves model output (Guo et al., 2024b). Some studies evaluate existing generative models on their effectiveness in handling low-resource languages (LRLs) such as Vallader (Bendel and Jabou, 2024), Odia (Agarwal et al., 2023), Faroese (Scalvini and Debess, 2024), and Marathi (Saini and Vidhyarthi, 2023). Instruction tuning is another category within our review paper set that tackles the data disparity problem by employing a technical approach. This method involves providing clear instructions for language models that enable them to adapt to unseen tasks and LRLs. In addition to adding machine translation instructions, word alignments have also been used to enhance the fine-tuning process (Mao and Yu, 2024). The efficacy of instruction tuning has also been assessed for adapting generative large language models for LRLs (Toraman, 2024). We have determined that two papers enlist cross-lingual information sharing as a method to bolster the data available for LRL models. Both papers build a cross-lingual information system that is responsible for collecting information for languages with more available training data when the information does not exist in the desired target language (Sorokin et al., 2022; Agarwal et al., 2022).
3.3.4. Multilingual Models
The category for multilingual models encapsulates all papers in this review that reference creating generative language models for more than a single language. 20% of the papers in this review were categorised for their use of multilingual models to address the issue of data scarcity when building generative models for LRLs. A number of NMT models naturally build generative language models that generate translations for pairs of languages (Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Pham et al., 2021; Usui and Komiya, 2023; Wang et al., 2018; Mi et al., 2024; Gao et al., 2019; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Guo et al., 2024b; Holiel et al., 2023; Ahmadnia and Dorr, 2019; Liu, 2022; Yu et al., 2020; Phan-Vu et al., 2017). In addition, several papers focused on building generative language models for a broader range of related languages. Some examples include a generative language model for 10 Indic languages (Niyogi and Bhattacharya, 2024); a model trained on a variety of Uralic languages (Downey et al., 2024); a multilingual model trained on a mixture of 250 high-resource and low-resource languages (Chang et al., 2023); and IndoNLG, which serves as a benchmark for generative tasks involving Indonesian languages (Cahyawijaya et al., 2021). Other notable multilingual models include a model paired with textbook-based prompts to teach LRLs to an existing language model (Guo et al., 2024b); a neural approach for next-word prediction across languages (Gerz et al., 2018); and a model that leverages cross-lingual information to extract content from high-resource languages like English and present it in another language (Sorokin et al., 2022; Agarwal et al., 2022). Additional work includes a translation model trained on a family of related Indonesian languages (Wongso et al., 2023), as well as a pre-trained language model designed for Chinese minority languages, such as Mongolian (Li et al., 2022a).
3.3.5. Monolingual Models
Papers tagged with the monolingual model category were almost as common as multilingual models in our systematic review, with the approach being adopted by 11% of the papers in this review. These papers encapsulate all studies that trained a generative language model for a particular language. This category differs from monolingual text data as some approaches involve building models for a single LRL without enlisting any form of data augmentation. Some of these approaches include training a Turkish LLM via continual training on various models that were not explicitly trained on Turkish data, such as Mistral-7B (Acikgoz et al., 2024); developing a Tibetan QA corpus generation model that generates text and applies BERT for grammar correction (Sun et al., 2019); building a sequence-to-sequence model for Slovenian as an LRL (Ulčar and Robnik-Šikonja, 2023); creating an Indonesian religious QA model (Rizqullah et al., 2023); and training a Bangla-based generative pre-trained transformer (Salim et al., 2023).
3.3.6. Adaptive Learning
Adaptive learning is featured as a common method for overcoming data scarcity, appearing in 11% of the papers examined for this systematic review. This method relates to taking pre-existing language models and resuming the training phase of the model to expose it to more data. LLamaTurk is based on the open source LLaMa-7b base and further trained on Turkish text data (Toraman, 2024). A similar approach is done for Turkish using Mistral-7b and GPT-2 as a base model (Acikgoz et al., 2024). Pre-trained models have also been fine-tuned to exploit cross-lingual capabilities for the Basque language (Otegi et al., 2020).
3.3.7. Family of Low-Resource Languages
There are a handful (9%) of papers that propose the idea of training generative language models on families of LRLs as opposed to broad multilingual models with a diverse range of languages. Therefore this study includes this as a separate category. Some papers that build a language model trained on multiple related languages point out that a model built for a family of languages can perform better than monolingual models for those languages (Downey et al., 2024; Niyogi and Bhattacharya, 2024; Cahyawijaya et al., 2021; Wongso et al., 2023).
3.3.8. Mass Translation
A pair (4%) of papers were found to perform mass translation as a means to bridge the data gap when modelling LRLs. Tools like Google Translate API and GPT 3.5/4 have been used to generate parallel datasets for training QA models for Bengali (Shafayat et al., 2024). One survey paper references mass translation as a method for translating missing values within a parallel corpus (Haque et al., 2021).
3.3.9. Miscellaneous
A variety of papers have unique approaches within this review set that have been categorised on their own for tackling the issue of data scarcity when building generative language models. One of these approaches pertains to extending the vocabulary of a large language model by merging the vocabulary of an existing model with the vocabulary of an LRL such as Turkish (Toraman, 2024). Another study adopts the concept of code-mixing, whereby a sequence-to-sequence model is trained on parallel data to create sentences comprising a mix of languages (Bhowmick et al., 2023).
Figure 7. Distribution of the technical methods used to overcome data scarcity in building generative language models by frequency (RQ2). Note that these categories are not mutually exclusive, as multiple methods may be employed in a single study.
Table 3. Overview of techniques used to address data scarcity in low-resource language settings (RQ2), mapped to the corresponding papers that employed or proposed them.
Technique Study
Augment Monolingual Data (Sun et al., 2019; Usui and Komiya, 2023; Gao et al., 2019; Marie and Fujita, 2020; Maimaiti et al., 2021; Guo et al., 2021a; Wang et al., 2018; Gao et al., 2020; Shafayat et al., 2024; Wu and Wang, 2023; Wang et al., 2021; Haque et al., 2021; Shi et al., 2022; Zhang et al., 2024b)
Back-Translation (Zhang et al., 2020; Marie and Fujita, 2020; Ahmadnia and Dorr, 2019; Yirmibeşoğlu and Güngör, 2023; Dandapat and Federmann, 2018; Wang et al., 2018; Pham et al., 2021; Hong et al., 2024; Berckmann and Hiziroglu, 2020; Pham Van and Le Thanh, 2022; Nissanka et al., 2020; Wang et al., 2022; Shi et al., 2022)
Prompt Engineering (Li et al., 2024b; Sorokin et al., 2022; Agarwal et al., 2022; Scalvini and Debess, 2024; Guo et al., 2024b; Bendel and Jabou, 2024; Mao and Yu, 2024; Toraman, 2024; Zhang et al., 2024a; Saini and Vidhyarthi, 2023; Agarwal et al., 2023)
Multilingual Models (Bhowmick et al., 2023; Cahyawijaya et al., 2021; Wongso et al., 2023; Niyogi and Bhattacharya, 2024; Downey et al., 2024; Chang et al., 2023; Tanwar and Majumder, 2020; Gerz et al., 2018; Liu, 2022; Babaali et al., 2024; Li et al., 2022a)
Monolingual Model (Sun et al., 2019; Ulčar and Robnik-Šikonja, 2023; Mi et al., 2024; Acikgoz et al., 2024; Salim et al., 2023; Rizqullah et al., 2023)
Adaptive Learning (Otegi et al., 2020; Yu et al., 2020; Holiel et al., 2023; Toraman, 2024; Acikgoz et al., 2024; Rizqullah et al., 2023)
Family of LRLs (Cahyawijaya et al., 2021; Wongso et al., 2023; Niyogi and Bhattacharya, 2024; Downey et al., 2024; Li et al., 2022a)
Mass Translation (Shafayat et al., 2024; Haque et al., 2021)
Code Mixing (Bhowmick et al., 2023)
Vocabulary Extension (Toraman, 2024)
3.4. RQ3: What Monolingual Text Data Augmentation Techniques Have Been Used to Build Generative Language Models for Low-Resource Languages?
There are a variety of technical methods that have been adopted for addressing and minimising the issue of limited training data for LRLs that have been explored as part of RQ2. Given the utility of data augmentation in computer vision tasks (Shorten and Khoshgoftaar, 2019), this systematic review dives deeper into applying similar synthesis processes for language modelling. Data augmentation techniques in this review relate to the papers that discuss synthesising new data from existing data for LRLs. We apply categories to each paper based on the type of text data augmentation used. The results of this section are displayed in Figure 8. The categories are organised into three broader groups based on the nature of the augmentation strategy: paraphrase-based techniques, grammatical transformation, and data enrichment. The following sections discuss these groups in order of frequency, starting with the most commonly used approaches.
3.4.1. Paraphrasing
The creation of new monolingual sentences through paraphrasing is the most popular approach for augmenting textual training data for building generative language models for LRLs. This paraphrasing process itself differs from paper to paper. These differing approaches include building a model that can use a set of random terms and subsequently produce QA pairs (Sun et al., 2019); enacting a variety of random word replacement strategies, such as rare-term replacement, simulated multiple reference training and soft context augmentation (Haque et al., 2021; Gao et al., 2019; Shi et al., 2022; Wang et al., 2018; Maimaiti et al., 2021); and making part-of-speech (POS) informed word replacements, whereby specific word classes are replaced instead of random replacement (Maimaiti et al., 2021).
3.4.2. Grammatical Transformation
Grammar manipulation is used in a variety of papers in this systematic review to create synthetic sentences by applying various syntactic and structural changes to original training sentences. Some of these grammar manipulation processes include using a reverse transcription grammar model to generate alternative sentences for a given input sentence (Wu and Wang, 2023) and introducing noise into existing training data to make it reflect user-generated text (Marie and Fujita, 2020).
3.4.3. Data Enrichment
One approach to data augmentation involved appending additional information to existing data. This information was in the form of metadata that encapsulated the title, access date, year and authors of a particular document (Usui and Komiya, 2023).
Figure 8. Distribution of the monolingual text data augmentation techniques identified in this systematic review by frequency (RQ3).
3.5. RQ4: Which Sources Are Publishing in the Area of Low-Resource Language Modelling?
There are a variety of sources publishing in the area of LRL modelling. Figure 9 indicates that the Association for Computational Linguistics (ACL) is the most active publisher in this area with 35.2% of the papers belonging to their anthology (Sorokin et al., 2022; Agarwal et al., 2022; Gerz et al., 2018; Dandapat and Federmann, 2018; Cahyawijaya et al., 2021; Mao and Yu, 2024; Gao et al., 2020; Marie and Fujita, 2020; Usui and Komiya, 2023; Wang et al., 2018; Gao et al., 2019; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Liu, 2022; Guo et al., 2024b; Wang et al., 2022; Otegi et al., 2020; Scalvini and Debess, 2024). ACL represents 19 different publications spanning 13 conference proceedings, three workshops, and three journals within this review. The Conference on Empirical Methods in Natural Language Processing (Zhang et al., 2020; Cahyawijaya et al., 2021; Wang et al., 2018) and the Transaction of the Association for Computational Linguistics (Marie and Fujita, 2020; Gerz et al., 2018; Wang et al., 2022) are the most frequently cited sources by ACL, both consisting of three papers in this review. Furthermore, the International Conference on Language Resources and Evaluation (Scalvini and Debess, 2024; Otegi et al., 2020) and the Annual Meeting of the Association for Computational Linguistics (Gao et al., 2019, 2020) are other conference papers by ACL that feature twice in this review. ACL also produces individual papers from nine other sources such as the Workshop on Multilingual Information Access (Agarwal et al., 2022); the Workshop on Technologies for Machine Translation of Low-resource Languages (Mao and Yu, 2024); the Conference on Machine Translation (Berckmann and Hiziroglu, 2020); the North American Chapter of the Association for Computational Linguistics (Sorokin et al., 2022); the Workshop on NLP for Similar Languages (Liu, 2022); the International Conference on Computational Linguistics (Guo et al., 2024b); the International Conference on Natural Language Processing for Digital Humanities (Usui and Komiya, 2023); the Annual Conference of the European Association for Machine Translation (Dandapat and Federmann, 2018); the European Chapter of the ACL (Li et al., 2024b).
The Association for Computing Machinery (ACM) is the joint-second most frequently cited publisher in this field, accounting for 18.5% of the papers included in this review (Bhowmick et al., 2023; Yirmibeşoğlu and Güngör, 2023; Maimaiti et al., 2021; Tanwar and Majumder, 2020; Pham Van and Le Thanh, 2022; Wu and Wang, 2023; Mi et al., 2024; Shi et al., 2022; Zhang et al., 2024b; Yu et al., 2020). Although ACM is the joint-second most frequently cited publisher among the sources from this review, its journal publication Transactions on Asian and Low-resource Language Information Processing is the largest individual source of papers in this review, as seen in Figure 10 (Bhowmick et al., 2023; Yirmibeşoğlu and Güngör, 2023; Maimaiti et al., 2021; Tanwar and Majumder, 2020; Mi et al., 2024; Shi et al., 2022; Zhang et al., 2024b; Yu et al., 2020). ACM is also responsible for a single paper via the Symposium on Information and Communication Technology (Pham Van and Le Thanh, 2022) and a single paper with the Asia Pacific Information Technology Conference (Wu and Wang, 2023).
Preprint papers from arXiv also constitute 18.5% of the total dataset (Pham et al., 2021; Niyogi and Bhattacharya, 2024; Downey et al., 2024; Hong et al., 2024; Chang et al., 2023; Shafayat et al., 2024; Wang et al., 2021; Toraman, 2024; Zhang et al., 2024a; Acikgoz et al., 2024). There are no further aggregations possible for arXiv papers given that the platform allows users to upload and share literature without being published by traditional academic journals or undergoing formal peer review. Additionally, IEEE is the publisher of 14.8% of the papers (Sun et al., 2019; Wongso et al., 2023; Nissanka et al., 2020; Salim et al., 2023; Saini and Vidhyarthi, 2023; Holiel et al., 2023; Rizqullah et al., 2023; Agarwal et al., 2023). IEEE Access is the only publication by IEEE that is cited more than once in this systematic review (Sun et al., 2019; Wongso et al., 2023). However, IEEE is the source of six other publications: the International Conference on Information and Communication Technology for Sustainable Development (Salim et al., 2023); the International Conference on Computer Systems and Applications (Holiel et al., 2023); the International Conference on Advances in ICT for Emerging Regions (Nissanka et al., 2020); the International Conference on Circuits (Agarwal et al., 2023); the International Conference of Advanced Informatics: Concept, Theory and Application (Rizqullah et al., 2023); the International Conference on Futuristic Technologies (Saini and Vidhyarthi, 2023). Moreover, Springer is the source of four papers consisting of 7.4% of the papers (Bendel and Jabou, 2024; Babaali et al., 2024; Li et al., 2022a; Haque et al., 2021). Springer is the publisher of Machine Translation (Li et al., 2022a; Haque et al., 2021) and the International Journal of Information Technology (Bendel and Jabou, 2024; Babaali et al., 2024), a pair of journals that are both cited twice in this review. Finally, Frontiers (Ulčar and Robnik-Šikonja, 2023), IJCAI (Guo et al., 2021a) and De Gruyter (Ahmadnia and Dorr, 2019) are all credited as the publishers of a single paper respectively, bringing each of their contributions to 1.9%. These publications include papers from: Frontiers in Artificial Intelligence (Ulčar and Robnik-Šikonja, 2023); the International Joint Conference on Artificial Intelligence (Guo et al., 2021a); and Open Computer Science (Ahmadnia and Dorr, 2019).
The location of the research institution affiliated with the first author of each paper is represented in Figure 12. China leads with the most publications, with the USA and India following behind. Figure 4 indicates the CORE ranking distribution for conference papers and the Journal Impact Factor 2024 distribution for journal papers. 33.3% of papers were not associated with a CORE-recognised conference. However, of the papers that are linked to a CORE-recognised event, 28.6% of them were ranked A*. On the other hand, the majority (57.9%) of journal publications were from Q3 journals. Figure 11 indicates that the number of empirical publications in this area has been increasing in recent years.
Figure 9. Distribution of publishers by frequency (RQ4).
Figure 10. Distribution of sources by frequency (RQ4).
Figure 11. Trend of survey vs empirical papers by year, and distribution of paper types by publication venue (e.g., conference, journal, symposium, etc.).
Figure 12. Distribution of publications by country.
3.6. RQ5: What Architectures Are Being Used to Produce Low-Resource Language Models?
A diverse set of architectures have been enlisted for LRL generation tasks. For this research question, we have assigned categories to each paper in this review depending on both the general architecture of the models that are built and the specific models that are used. Figure 13 indicates that the transformer architecture dominates as the architecture of choice for 76% of the papers in this review. In terms of the papers that model LRLs, the majority of them employ a transformer-based approach. Recurrent Neural Network-based implementations are the second most frequently adopted architecture for LRL modelling, with nine studies in this systematic review opting to utilise them. The Long Short-Term Memory (LSTM) categorisation applies to six papers. Given that LSTM-based approaches are inherently RNNs, it is worth highlighting that the categorisations are not mutually exclusive. 4% of papers in this review enlist Generative Adversarial Networks (GAN) for building LRL generation models. Additionally, a Gated Recurrent Unit (GRU) approach is implemented by one study, while another study employs a Statistical Machine Translation (SMT) architecture. In terms of the specific models being used to produce LRL models, Figure 14 showcases the most frequently implemented models. GPT-2 appears most often in the studies with 11% of papers enlisting it for LRL modelling. XLM, MT5 and mBERT are each modelled by 7% of studies. T5 is used by 6% papers for LRL modelling. Furthermore, mBART and Mistral-7b are both enlisted by 4% of papers. The rest of the models are used once and include variations of LLama, BERT and BART models.
Figure 13. Distribution of the model architectures used by papers in this systematic review by frequency (RQ5).
Figure 14. Model distribution by frequency within the systematic review (RQ5). Note: some studies report general architectures rather than specific models.
3.7. RQ6: How Effective Have Methods for Overcoming Data Scarcity Been at Improving Performance for Different NLG Tasks?
Text data augmentations have been used by a variety of studies in this review to improve the performance of language generation tasks for LRLs. Back-translation has been found to improve BLEU scores for translation tasks in comparison with no augmentation (Pham et al., 2021; Hong et al., 2024; Berckmann and Hiziroglu, 2020; Yirmibeşoğlu and Güngör, 2023; Nissanka et al., 2020; Dandapat and Federmann, 2018; Pham Van and Le Thanh, 2022). This contrasts slightly with a minor BLEU improvement found by another study that enlists back-translation to improve Cherokee-English machine translation (Zhang et al., 2020). However, the study attributes the limited performance improvement to the lack of in-domain training data. Using a combination of synthetic code-mixed and mixed-script with original data has been found to improve BLEU, METEOR and ROUGE scores for Indic language translation (Bhowmick et al., 2023). Synthesising new text data via paraphrasing as a data augmentation technique has also been shown to increase the BLEU performance of translating from an LRL to English (Guo et al., 2021a). Paraphrasing has improved BLEU results at both a character and word level for question-answering corpus generation for Tibetan (Sun et al., 2019). Paraphrasing has also been used to improve BLEU scores by boosting data in low-resource tasks such as dialogue generation (Gao et al., 2020). Reordering monolingual sentences as a form of text data augmentation has also yielded BLEU score improvements across a baseline of other data augmentation techniques (Wu and Wang, 2023). Similarly, synthesising user-generated text has been used as an augmentation technique to improve BLEU scores for NMT (Marie and Fujita, 2020). Training a language model on a family of related languages is an effective data augmentation technique whereby a study has built an Indonesian language model that achieves similar results to state-of-the-art models with only  20% of the number of parameters (Cahyawijaya et al., 2021). Another approach likens multilingual language modelling of related languages to a 33% increase in the size of the monolingual dataset (Chang et al., 2023). Mass translation as a text data augmentation technique has also been found to improve COPA results for Bengali (Shafayat et al., 2024). Word-level augmentations have been successful at improving BLEU scores across other augmentation techniques for low-resource NMT (Maimaiti et al., 2021; Wang et al., 2018; Gao et al., 2019).
3.8. RQ7: How Scarce Is the Data in Studies Tackling Low-Resource Language Modelling?
As part of this review, we determine the quantity of data that has been used to model generative language models for LRLs. This process involved investigating and exploring the corpora that have been used to train models for LRL in this review. Given the ambiguity surrounding corpus statistics in many of the papers reviewed, this research question focuses on whether the study specifies the number of sentences used to train the LRL model. We found that three different datasets were used to train generative language models for Khmer (Agarwal et al., 2022; Mao and Yu, 2024; Pham Van and Le Thanh, 2022) with data ranging from 70k-579k sentences. Figure 15 highlights the languages with a single data source, where no additional sources are available. This visualisation indicates that Minnan has been modelled with the largest sentence count of 1.2m sentences (Gerz et al., 2018). From this aggregation, a count of 390k sentences is the average count for training language models for these LRLs. In several studies, data is quantified in terms of size in gigabytes rather than sentence count. For instance, Bengali models were trained on 26GB of data (Salim et al., 2023), while Estonian, Turkish, and Indonesian models used 6.1GB (Downey et al., 2024), 180GB (Acikgoz et al., 2024), and 23.8GB (Cahyawijaya et al., 2021) of data, respectively.
Figure 15. Distribution of sentence counts by language (RQ7).
3.9. RQ8: How Are the Models in These Studies Being Evaluated?
Numerous evaluation methods have been enlisted to assess the performance of NLG models for LRLs in this review. Figure 16 indicates that BLEU dominates as the evaluation method of choice with 61% of papers using it to gauge the efficacy of their models (Salim et al., 2023; Bhowmick et al., 2023; Tanwar and Majumder, 2020; Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Cahyawijaya et al., 2021; Mao and Yu, 2024; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Pham et al., 2021; Wu and Wang, 2023; Sun et al., 2019; Guo et al., 2021a; Gao et al., 2020; Marie and Fujita, 2020; Usui and Komiya, 2023; Wang et al., 2018; Mi et al., 2024; Gao et al., 2019; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Holiel et al., 2023; Ahmadnia and Dorr, 2019; Yu et al., 2020; Phan-Vu et al., 2017; Li et al., 2022a; Otegi et al., 2020; Zhang et al., 2024a). ROUGE is used by 9% of studies (Bhowmick et al., 2023; Saini and Vidhyarthi, 2023; Holiel et al., 2023; Li et al., 2022a), hence it is the second most frequently used evaluation method in this review. Perplexity, chrF, and METEOR are each used in 7% of the papers to evaluate their models (Salim et al., 2023; Chang et al., 2023; Bhowmick et al., 2023; Maimaiti et al., 2021; Toraman, 2024; Saini and Vidhyarthi, 2023; Holiel et al., 2023; Li et al., 2022a; Scalvini and Debess, 2024; Zhang et al., 2024a; Vo and Tran, 2023). Furthermore, five evaluation methods each appear in 6% of the studies. These methods include human evaluation (Niyogi and Bhattacharya, 2024; Dandapat and Federmann, 2018; Hong et al., 2024); chrF++ (Mao and Yu, 2024; Li et al., 2024b; Guo et al., 2024b), sacreBLEU (Wongso et al., 2023; Hong et al., 2024; Liu, 2022); COMET (Mao and Yu, 2024; Hong et al., 2024; Guo et al., 2024b); and F1-score (Sorokin et al., 2022; Rizqullah et al., 2023; Otegi et al., 2020).
Figure 16. Distribution of evaluation method by frequency (RQ8).
3.10. RQ9: What Is the NLG Task?
A variety of generative models have been created for LRLs by studies in this review. Figure 17 showcases each NLG task by frequency. Given that some papers attempt to build models for multiple generative tasks, the categorisations are not mutually exclusive. Translation is the most popular generative task category in this review, with 57% of the studies focusing on the task (Bhowmick et al., 2023; Agarwal et al., 2022; Gerz et al., 2018; Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Toraman, 2024; Acikgoz et al., 2024; Yirmibeşoğlu and Güngör, 2023; Cahyawijaya et al., 2021; Wongso et al., 2023; Mao and Yu, 2024; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Vo and Tran, 2023; Downey et al., 2024; Pham et al., 2021; Haque et al., 2021; Shi et al., 2022; Wu and Wang, 2023; Sun et al., 2019; Guo et al., 2021a; Wang et al., 2018; Li et al., 2024b; Yu et al., 2020; Phan-Vu et al., 2017; Guo et al., 2024b; Wang et al., 2022; Otegi et al., 2020; Zhang et al., 2024a). Furthermore, general language modelling is the second most popular language generation task in this review with 24% of the papers tackling it (Niyogi and Bhattacharya, 2024; Salim et al., 2023; Chang et al., 2023; Gerz et al., 2018; Toraman, 2024; Acikgoz et al., 2024; Cahyawijaya et al., 2021; Downey et al., 2024; Berckmann and Hiziroglu, 2020; Holiel et al., 2023; Li et al., 2022a; Ulčar and Robnik-Šikonja, 2023). We categorise general language modelling as research relating to building large pretrained models as a base for various downstream tasks. Question-answering is also somewhat popular with 15% of the papers building models for the task (Shafayat et al., 2024; Sorokin et al., 2022; Agarwal et al., 2022; Sun et al., 2019; Rizqullah et al., 2023; Otegi et al., 2020; Vo and Tran, 2023). Summarisation (Gao et al., 2020) and dialogue generation models are also modelled by a single paper in this systematic review (Holiel et al., 2023).
Figure 17. Distribution of language generation task by method (RQ9).
4. Discussion
This section discusses the findings of the included studies, obtained through the data collection process, concerning the research sub-questions that inform the overall research objective of this systematic review. Each research sub-question is discussed in a separate subsection.
4.1. RQ1: What Low-Resource Languages Have Already Been Modelled by Other Studies?
Many LRLs have been modelled by studies in this review for a variety of different natural language generation tasks. Given that Indo-European languages dominate the literature as the most frequently modelled LRLs with 133 different Indo-European languages being modelled, this could suggest that even among LRLs there is a disparity in terms of their representation. While many LRLs being modelled belong to the Indo-European language family, there are some language families such as the Aymaran and the Tupian language families whereby only a single language was modelled from their language family. It is worth considering that this disparity could be attributed to region-specific language families lacking the necessary resources to produce research in this area. This disparity indicates that there is a need for more research that is representative of all language families within the LRL group. Although this review is focused on bridging the data gap for modelling LRLs, there are other barriers such as research demand and sourcing expertise in the area. This could indicate that categorising the resources of a language requires considering the research interest, in addition to considering the amount of training data available for such an LRL.
4.2. RQ2: What Technical Methods Have Been Used to Overcome Data Scarcity Building Generative Language Models?
A wide range of technical methods have been used to overcome the challenge of data scarcity when building generative language models for LRLs. Similar to the results section for RQ2, the following discussion is divided by technical methods used to overcome data scarcity in order of frequency, starting with the most commonly used approaches.
4.2.1. Monolingual Text Data Augmentation
The process of synthesising new data from monolingual text data was the most frequently used approach for overcoming data scarcity. This often involved performing transformations on existing sentences. Some of these synthesis methods relate to replacing low-frequency terms with other terms chosen by an LSTM model (Fadaee et al., 2017; Haque et al., 2021; Shi et al., 2022). These approaches are bounded by the performance of the LSTM model chosen for replacement. Errors in the generated text produced by the chosen LSTM model could propagate to future models. In the context of language generation tasks such as question-answering and paraphrase generation, rare words could be indicative of domain-specific information or carry contextual information that would be lost with an LSTM replacement. This could contribute to nonsensical or vague information being produced by the models. Similarly, instead of performing a replacement on low-frequency terms, other studies have exploited POS tagging (Maimaiti et al., 2021) to inform the replacement process. While this method of replacing specific POS tags with similar terms using a paraphrase table could work for some LRLs, it could be challenging to apply this to LRLs with a complex morphological language such as Gaeilge, whereby altering a term in a sentence could corrupt the grammatical agreement of the sentence (Legate, 1999). This POS-informed replacement could be improved by adding a final neural approach for grammar correction. Moreover, another approach involves randomly replacing a word with a similar meaning using a pre-trained language model (Haque et al., 2021; Gao et al., 2019). This method could be enhanced by combining it with the previously mentioned POS-informed replacement. Performing contextualised replacements on specific POS tags could circumvent the issue of agreement and other grammatical errors. Replacements have also been done randomly without using a language model (Haque et al., 2021; Shi et al., 2022; Wang et al., 2018). Yet, without context-aware replacement strategies, randomly swapping terms could corrupt the semantics of the text.
Alternatively, some studies duplicate text in a corpus and use a model to paraphrase them (Shi et al., 2022; Guo et al., 2021a; Gao et al., 2020). This could be an effective way of bolstering a dataset for LRLs, as synthesising new sentences that conserve the original meaning of the sentence would expose the language model to more diverse and representative data. A related approach involves adopting grammar correction to improve generated text for question-answering in Tibetan (Sun et al., 2019). This BERT-based strategy for fixing grammatical errors in output sentences could help to reduce erroneous text generation for technical methods that involve randomly replacing terms or performing contextually unaware sentence restructuring. It could be an additional step to ensure that the grammar of generated text is not corrupted. However, it is worth highlighting that this would hinge on the availability and efficacy of grammar correction tools and models. Another study performs corpus expansion by using a reverse transcription model to generate new sentences through grammatical reordering of the original sentences (Wu and Wang, 2023). Given that some languages are more flexible in terms of their word order (Dryer, 2007), enlisting a model to generate grammatically correct alternative sentences could enrich a dataset of LRL text by increasing the diversity of language. Other studies have synthesised data by reordering target language text to align with the structure of a source language (Haque et al., 2021). While this is a beneficial approach for machine translation tasks, this approach could produce structurally incorrect sentences that do not conform to the morphological features of the language. In terms of a dialogue model, training on text data with instances of poor grammar could create a model that is not representative of the language itself. One approach for augmenting and synthesising parallel data involves performing mass translation of text data from a source language to a target language using tools such as GPT 3.5 and Google Translate API (Shafayat et al., 2024).
Another approach uses translation tools to impute missing text examples in parallel datasets (Haque et al., 2021). While these papers use translation tools to enhance translation models, the same approach could be adapted to translate textual data from a higher-resourced language like Chinese to a target LRL. However, the quality of the synthesised data hinges on the existence and availability of performant translation models. Building generative language models on translated text could propagate the errors from translation models through to generative language models. One study references building user-generated text for NMT by employing noise and stylistic changes to textual data (Dandapat and Federmann, 2018). While this method produces synthetic data, it could be more suited for the specific task of translating user-generated content, as opposed to augmenting data for other text generation tasks such as question-answering, whereby noise could distort information.
4.2.2. Back-Translation
Back–translation is the second most frequently used text data augmentation technique in this review for tackling data scarcity when building generative text models for LRLs (Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Pham et al., 2021; Shi et al., 2022; Marie and Fujita, 2020; Wang et al., 2018; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Wang et al., 2022). While round-trip back-translation can bolster monolingual data (Guo et al., 2021a), back-translation can be a beneficial tool for building synthetic parallel datasets for translation tasks. However, back-translation suffers from similar pitfalls as any other translation-based text data augmentation such as mass translation. The efficacy of back-translation is heavily dependent on the quality and performance of the translation tools that are being used to synthesise these back-translated inputs. While back-translation has improved translation performance for modelling a variety of LRLs (Berckmann and Hiziroglu, 2020; Pham Van and Le Thanh, 2022; Nissanka et al., 2020), this process could consequently produce synthetic text data that disrupts the semantics and grammar of the original text data. Further investigations into boosting LRL data for generative language modelling through the use of back-translation could be an insightful experiment.
4.2.3. Multilingual Models
Building multilingual models is an approach adopted by many of the papers in this study to address the lack of training data for LRLs. Given that NMT models require a mapping from a source language to a target language and vice versa, there are many NMT models discussed in this review (Maimaiti et al., 2021; Nissanka et al., 2020; Dandapat and Federmann, 2018; Yirmibeşoğlu and Güngör, 2023; Pham Van and Le Thanh, 2022; Saini and Vidhyarthi, 2023; Pham et al., 2021; Usui and Komiya, 2023; Wang et al., 2018; Mi et al., 2024; Gao et al., 2019; Hong et al., 2024; Zhang et al., 2020; Berckmann and Hiziroglu, 2020; Li et al., 2024b; Guo et al., 2024b; Holiel et al., 2023; Ahmadnia and Dorr, 2019; Liu, 2022; Yu et al., 2020; Phan-Vu et al., 2017). Cross-lingual transfer learning can be leveraged for various generation tasks by building models for a variety of LRLs (Niyogi and Bhattacharya, 2024; Chang et al., 2023; Downey et al., 2024; Cahyawijaya et al., 2021; Guo et al., 2024b; Sorokin et al., 2022; Gerz et al., 2018; Agarwal et al., 2022; Wongso et al., 2023; Li et al., 2022a). Although one study trains multilingual models from a mixture of 250 high and LRLs (Cahyawijaya et al., 2021), most of the studies create models for related languages (Niyogi and Bhattacharya, 2024; Downey et al., 2024; Cahyawijaya et al., 2021; Wongso et al., 2023; Li et al., 2022a). The study that experiments with 250 high and low resource-languages (Cahyawijaya et al., 2021) argues that building multilingual models can be likened to augmenting the training dataset by 33%. The study explores the potential benefits of training similar or maximally related languages together, suggesting that this approach could enhance performance. These findings present an alternative to monolingual text data augmentation strategies for addressing data scarcity. However, the results also indicate that some multilingual language models can underperform when compared to their monolingual counterparts, highlighting the importance of selecting compatible language groups for multilingual modelling. Alternatively, a pair of studies train models on multiple languages to exploit cross-lingual information (Sorokin et al., 2022; Agarwal et al., 2022). Both of these approaches involve collating information from more resourced languages to generate more relevant and accurate text in a target language. However, this process hinges on such relevant information existing in other languages. Monolingual text data augmentations cannot synthesise new information, therefore this cross-lingual approach could expose LRL models to more diverse information. For example, some LRLs may lack news services that provide coverage of global affairs in their language. Consequently, facilitating cross-lingual information sharing can help bridge this gap, ensuring that speakers of these languages have access to a broader range of knowledge not written explicitly in the target language, such as global news. Another study investigates training a sequence-to-sequence model using parallel data to generate sentences consisting of a mix of both languages (Bhowmick et al., 2023). While this could create seemingly nonsensical data, code-mixing could accurately reflect some communities such as Irish people, as they often inject English into their Irish conversations and vice versa (Laoire, 2016). Therefore, code-mixing could be a valid method of augmenting data for a pair of languages that are frequently used in conjunction with each other.
4.2.4. Prompt Engineering
Several papers in this study attempt prompt engineering techniques on existing models to investigate their capacity to adapt to different LRLs (Li et al., 2024b; Zhang et al., 2024a; Guo et al., 2024b; Bendel and Jabou, 2024; Agarwal et al., 2023; Saini and Vidhyarthi, 2023; Scalvini and Debess, 2024). An interesting approach enlists a programmer-interpreter process to continuously refine the output of a language model by using smaller models to perform corrections and further changes (Li et al., 2024b). While this process could be beneficial for synthesising additional training data for LRLs with a lack of training data, it can produce nonsensical text that does not conform with the grammatical and semantic features of the target language. Two papers in this review enrich generative language prompts by injecting additional information into the prompt itself (Guo et al., 2024b; Zhang et al., 2024a). One of these approaches involves the Talent framework whereby a textbook is created for an LRL (Guo et al., 2024b). This textbook is subsequently utilised to guide large language models to translate by including vocabulary and language examples with the prompt. While this approach reflects how humans learn a new language, it hinges on the multilingualism of the large language model. Similarly, another study uses a few-shot approach that appends example translations and definitions to guide the translation model (Zhang et al., 2024a). If these models being prompted and guided do not have cross-lingual knowledge or a broad vocabulary range that can represent LRLs, they could struggle to adapt to a language regardless of additional information and context in the prompt. There are several studies in this review that simply assess the capability of existing pre-trained generative language models for a variety of LRLs (Bendel and Jabou, 2024; Agarwal et al., 2023; Saini and Vidhyarthi, 2023; Scalvini and Debess, 2024). A study evaluating GPT-SW3 for Faroese makes the case for using a multilingual language model for a family of languages to augment data for a lower resource language within the family (Scalvini and Debess, 2024). This is an interesting perspective that warrants experimentation. Further research is required to investigate the efficacy of training a language model for an LRL by using such a dataset augmented with text generated by a multilingual model from the same language family.
4.2.5. Monolingual Models
In contrast with building multilingual models, some studies assess whether existing monolingual data is sufficient for modelling LRLs or if the scarcity of such data prevents effective modelling of LRLs (Acikgoz et al., 2024; Sun et al., 2019; Ulčar and Robnik-Šikonja, 2023; Rizqullah et al., 2023; Salim et al., 2023). An insightful approach involves enlisting transfer learning to adapt existing generative language models for a specific LRL. Using this approach, a Turkish language model was created by exposing existing models to additional training data (Acikgoz et al., 2024). Exploiting transfer learning allowed the model to leverage knowledge from the original languages it was trained on, to better understand and generate text in Turkish, even with limited Turkish-specific data. The paper found that adapting the Mistral model for Turkish improved accuracy across reasoning and question-answering tasks, indicating that adaptive learning can improve LRL modelling. On the other hand, one paper further trains various existing generative models on ChatGPT-generated QA pairs for the Indonesian language (Rizqullah et al., 2023). However, their use of text data sourced from ChatGPT biases their model. Their model is bounded by ChatGPT’s capacity to generate text in the Indonesian language. Similarly, the language used by ChatGPT to create questions might not reflect questions that would be posed by humans. Despite these limitations, using data generated by an existing language model could serve as a foundation and baseline for building generative models for LRLs.
An alternative approach builds a QA corpus generation model for the Tibetan language (Sun et al., 2019). Although it employs BERT as a method to correct grammatical errors in the generated QA text, the paper references that the generation model is not aware of and does not adhere to the grammatical characteristics of the Tibetan language. Their approach randomly masks words in their dataset and uses BERT to replace the term. While this could generate valid grammatical corrections, the model could produce nonsensical or even unrelated terms that lose the context of the original term. These replacements could also corrupt the grammar of the sentences. This word-level approach does not account for wider sentence-level grammatical errors. Another approach involves training a sequence-to-sequence model for the Slovenian language (Ulčar and Robnik-Šikonja, 2023). Interestingly, the paper finds that their small T5 models outperform the multilingual models, yet the large multilingual model outperforms their large T5 model. This highlights that there is some evidence that monolingual models can outperform multilingual models for less-resourced languages. However, these results also suggest that some languages could achieve better performance by leveraging multilingualism and transfer learning. The Bangla language is modelled using a considerable amount of data without using any specific text data augmentation techniques (Salim et al., 2023). However, there are no baseline comparisons in this study. It would be insightful to compare the performance of the monolingual Bangla model with other multilingual models that have been trained on Bangla text.
4.2.6. Adaptive Learning
In terms of adaptive learning, two papers opted to extend the training of preexisting models by exposing them to Turkish text data (Toraman, 2024; Acikgoz et al., 2024). This interesting approach takes advantage of cross-lingual information through transfer learning and subsequently yields promising results. One of the papers highlights that continued training with raw text and that extending the vocabulary size did not have an impact on performance (Toraman, 2024). The other paper interestingly highlights that resuming the training process on multilingual models can trigger catastrophic forgetting of their primary language knowledge (Acikgoz et al., 2024). While exploiting transfer learning could improve performance for generative language models, the paper indicates that large models like Gemma7B still perform better without fine-tuning. This could indicate that better results can be obtained by collating a larger more representative dataset and building models with more parameters. Word alignments have also been added for instruction tuning to enhance the performance in translation tasks for LRLs, indicating that instruction fine-tuning could be an effective measure of overcoming data scarcity in LRL translation tasks (Mao and Yu, 2024).
4.2.7. Family of Low-Resource Languages
A variety of papers target building generative language models for families of related languages (Niyogi and Bhattacharya, 2024; Cahyawijaya et al., 2021; Wongso et al., 2023; Downey et al., 2024; Li et al., 2022a). One paper successfully builds multilingual language models for the Uralic language family (Downey et al., 2024). Although the study indicates that LRLs benefit from being trained as part of a wider Uralic multilingual language model, even the more resourced languages such as Hungarian and Russian also benefit. They suggest it is infeasible to model some LRLs such as Komi and Moksha without exploiting multilingualism and transfer learning. This chimes with the results of IndoNLG (Cahyawijaya et al., 2021), as their model achieves similar results to another large multilingual model (mBART-large), even though they use  20% of the number of parameters. One of the previously mentioned studies that experiments with 250 high and LRLs (Chang et al., 2023) suggests that building multilingual models can be compared to augmenting the dataset for a monolingual model by 33%. It could be both beneficial and insightful to investigate the impact of training multilingual models for closely related languages. It would be interesting to observe whether a model trained on a family of LRLs could be compared to an even larger increase in data. These results suggest that building generative language models for a family of LRLs could not only yield strong performance but also reduce the need to train separate models for each language, making the approach more cost-effective and environmentally sustainable.
4.2.8. Mass Translation
Mass translation has been used to generate parallel data by employing services like Google Translate API and GPT 3.5/4 (Shafayat et al., 2024). As mentioned previously, mass translation can be used for boosting monolingual data and also for augmenting parallel data. While this process has been executed to create QA systems for Bengali (Shafayat et al., 2024), the results are highly dependent on the translation quality of these translation services. Although similar methods are used to bolster LRL datasets for translation, this process could generate inaccurate text that could impede QA and other generative tasks such as dialogue generation. Another paper uses translation tools to impute missing data in parallel datasets (Haque et al., 2021). While this could be an effective way of dealing with missing values in a dataset, it could also corrupt the semantic information in text data which would result in inaccurate information being generated for tasks such as QA.
4.2.9. Miscellaneous
One unique method mentioned within this review pertains to extending the vocabulary size of existing generative models that have been predominantly trained on English textual data (Toraman, 2024). While this is an interesting approach, the paper finds that adapting existing models by extending the vocabulary does not have a significant impact on performance. However, it highlights that continual training can improve perplexity. This could indicate that additional training outweighs the vocabulary size when training generative language models for LRLs. In addition, one paper applies the concept of code-mixing as a method of augmenting text data (Li et al., 2024b). While this approach is previously discussed within the scope of being a multilingual approach, code-mixing should also be referenced with its capacity to capture more community-centric language variations. Although code-mixing can be used to build models that generate text in multiple languages, code-mixing in this way is not necessarily modelling each of these languages, but it is rather modelling the intersection of them. The paper suggests that this approach can create synthetic data with a Code-Mixing Index comparable with existing authentic datasets. This could be significant for building generative language models for LRLs that closely relate to another language. Given that English and Irish terms are often used interchangeably in Ireland (Laoire, 2016), a code-mixing approach could subsequently build a generative language model that more accurately captures the colloquial language of that community. Not only has code-mixing been shown to be effective in addressing data scarcity when building generative language models, but it can also enhance the representation of communities that naturally interweave multiple languages into their vernacular.
4.3. RQ3: What Monolingual Text Data Augmentation Techniques Have Been Used to Build Generative Language Models for Low-Resource Languages?
Given the prevalence of data augmentation in computer vision tasks, monolingual text augmentation techniques should be explored in further detail. Furthermore, this category encapsulates a discussion on the algorithms and processes presented in the results section for RQ3, that could be used to synthesise new training data for LRLs.
4.3.1. Paraphrasing
Paraphrasing was the most referenced data augmentation technique in our review set of papers. Several papers employ existing or additional models to paraphrase authentic text data and subsequently synthesise new sentences that preserve the semantic information of the original data (Maimaiti et al., 2021; Sun et al., 2019; Guo et al., 2021a; Gao et al., 2020). One paper creates paraphrased text by using randomly generated text sequences as input to a model trained on Tibetan QA pairs to generate new synthetic QA pairs (Sun et al., 2019). The study subsequently employs BERT to apply some word-level grammatical fixes to the output text. While the results of their generation model indicate that some text inputs can produce coherent QA pairs, these results can also ignore the grammar characteristics of the target language. Although this approach of utilising a language model to paraphrase monolingual text data can produce grammatical errors, this could be compared to the distortions that can occur with back-translation. Therefore, it could be insightful to enlist paraphrasing tools to bolster text data in the context of machine translation tasks. In contrast, this potentially erroneous and nonsensical output could have a detrimental effect on the training of models for other generation tasks like dialogue generation and abstractive text summarisation.
A variety of studies in this review produce paraphrased text by enacting different types of word replacement strategies (Haque et al., 2021; Gao et al., 2019; Shi et al., 2022; Wang et al., 2018; Maimaiti et al., 2021). Another approach refers to creating numerous duplications of original sentences and applying random replacements on the duplicated sentences to produce new paraphrased sentences (Guo et al., 2021a). Much like the word-level grammatical checks used to improve the quality of generated QA pairs (Sun et al., 2019), performing word-level augmentations by replacing terms with random or context-informed words could also distort the original interpretation of the sentence. One of these approaches tackles the shortcomings of word-level replacements by utilising POS tag-based augmentations (Maimaiti et al., 2021). This is an interesting process as choosing specific POS tags and replacing them could reduce the risk of grammatical errors produced through synthesising data. If a language has word classes that do not change based on agreement with other words in a text, a contextualised word replacement based on POS tags could produce more grammatically correct augmented sentences. For example, adverbs do not affect agreement in English sentences, therefore they could be replaced with semantically similar terms without corrupting the grammar of the sentence. However, it is worth highlighting that this approach is highly dependent on the quality of available POS tagging tools and models for a given LRL. Not only does the quality of POS-informed replacements hinge on the availability of POS tagging tools, the efficacy of the replacement strategy also directly impacts the quality of the paraphrased text. Although POS-based augmentations could reduce the risk of synthesising grammatically incorrect text, the replacement strategy, neural-based or not, could corrupt or dilute the original meaning of the text being augmented. While this POS-informed approach is used to improve NMT for LRLs, this data augmentation method could also bolster general language generation models, particularly for LRLs that consist of word classes which do not affect the grammar of other terms in a sentence.
4.3.2. Grammatical Transformation
There are a couple of papers that perform reordering on monolingual sentences as an approach for synthesising new data from existing text data (Wu and Wang, 2023; Marie and Fujita, 2020). While these methods can be viewed as a form of paraphrasing, they explicitly perform a degree of restructuring. These approaches involve duplicating existing sentences and reconstructing them as a means to augment the entire dataset. One paper uses a reverse transcription grammar model to generate diverse sentences for given input sentences (Wu and Wang, 2023). While this method can produce grammatically correct text, the reordering process could corrupt the original meaning of the sentence. Therefore, it is possible that this method could be beneficial for NMT tasks in the same way that back-translation improves translation performance. In terms of generating text for dialogue generation or abstract summarisation, training text data containing many distortions of factual content could produce poor models. However, this process of reordering a sentence could be beneficial for LRLs that have word order flexibility. Another approach relates to synthesising new text data by introducing noise to reformat input text data to resemble user-generated text (Marie and Fujita, 2020). This is achieved by applying various stylistic changes such as introducing spelling errors, slang and informal expressions. While this approach is applied to improving NMT systems for user-generated content, there could be applications of this method for building grammar alteration models. Models could be built for adapting or translating from one sentence structure to another within the same language. This would still follow the sequence-to-sequence nature of NMT but subsequently produce various augmentations of an original input sentence. Not only could producing such synthesised sentences improve NMT, but this form of text data augmentation could also empower other generative tasks like dialogue generation and QA tasks for LRLs.
4.3.3. Data Enrichment
The approach of enriching existing data by appending additional information to it is quite a straightforward approach that relates to adding the titles of books to each example sentence for building a translation model from Historical Japanese to Contemporary Japanese (Usui and Komiya, 2023). While the impact of this data enrichment technique is not explored in depth, more research is required to determine if adding metadata to text documents could help with text generation tasks. It could be worth exploring the impact of including other data such as the access date, the year, the authors and the title of the training data. Performing this process could help to ground the generative language model through the inclusion of relevant information for each text example.
4.4. RQ4: Which Sources Are Publishing in the Area of Low-Resource Language Modelling?
The results section for RQ4 indicates that there is a wide variety of sources publishing in the area of LRL modelling. Given the dominance of ACL publications in this review spanning journals, workshops and conferences alike, it is clear that it has a central influence on the dissemination of literature relating to modelling LRLs. The high frequency of publications from specific events such as EMNLP and LREC could suggest that these are key venues for LRL research. The prevalence of EMNLP papers could be attributed to the empirical nature of language modelling, and the frequency of LREC papers could be linked to the language resource element of building language models for LRLs. While ACM is the publisher of fewer events in comparison with other organisations such as ACL and IEEE, the Transactions on Asian and Low-Resource Language Information Processing journal is the most frequently sourced publication in this review. This could highlight that this journal is highly influential in terms of LRL modelling. It is also worth mentioning the prevalence of arXiv preprints in this review. While arXiv papers are not necessarily associated with traditional academic journals or undergo formal peer review, their prevalence could suggest that the field of LRL modelling is a rapidly evolving research topic. Given that arXiv papers consist of 19% of the review dataset, they could introduce limitations relating to the quality of research in the review process. However, we mitigate this effect by enlisting the quality assessment question relating to the rankings of conferences and journals. On the other hand, IEEE has more of a multi-disciplinary influence as it is the source of eight different publications relating to various fields within ICT and engineering. This could highlight that there are potential applications for LRL modelling in a variety of domains. Considering that Springer is the source of five papers in this review, this could indicate that there is demand for journals that relate to specialised fields to include literature that tackles LRL modelling. Single contributions and participation from other publishers such as Frontiers, IJCAI, and De Gruyter could indicate that the research area of LRL modelling could be rising in prevalence. Ultimately, there is a broad range of sources that span eight different publishers in the field of LRL modelling.
4.5. RQ5: What Architectures Are Being Used to Produce Low-Resource Language Models?
The prevalence of transformer-based models in this review reflects the transformative nature of the architecture for generative language modelling. Given the rise in popularity, adoption and performance of transformer-powered services such as ChatGPT, it is unsurprising that it dominates as the most frequently implemented model architecture in this review. Figure 13 indicates that transformer-based models could be the most suitable architecture for modelling LRLs. RNNs, and more specifically LSTMs are also somewhat popular architectures for modelling LRLs in this review. This could reflect the evolution of the state-of-the-art architecture over time as RNN/LSTM-based approaches would have been the gold standard before the transformer was introduced (Mienye et al., 2024). While GAN, GRU and SMT-based methods have been used by studies in this review, they are not as popular for modelling generative models for LRLs. Their unpopularity could suggest that they are not the optimal architectures for LRL generation. Interestingly, GPT-2 is the most popular model across all of the studies seen in this review. Given that GPT-2 is an open-source predecessor to the widely popular GPT-3 and GPT-4 models by OpenAI, its popularity could indicate that it is suitable for LRL modelling. The size of GPT-2 small could make it appealing to LRL researchers, as it does not require as much training time and computational power to build it as other prevalent models such as the Llama3 family of models (Radford et al., 2019; Dubey et al., 2024). XLM and sequence-to-sequence models such as mT5 and T5 are also frequently used for LRL modelling. Their popularity could be attributed to the considerable number of models in this review that relate to translation, as the task is naturally a sequence-to-sequence task. Ultimately, there is a wide variety of models being used to build generative models for LRLs, although transformer-based approaches are the most frequently used for this task.
4.6. RQ6: How Effective Have Methods for Overcoming Data Scarcity Been at Improving Performance for Different NLG Tasks?
While some text data augmentation techniques are effective at improving the performance of different NLG models, it is worth considering how such methods compare across different NLG tasks. Translation as a generative language task can be improved with a number of augmentation approaches such as back-translation (Pham et al., 2021; Hong et al., 2024; Berckmann and Hiziroglu, 2020; Yirmibeşoğlu and Güngör, 2023; Nissanka et al., 2020; Dandapat and Federmann, 2018; Pham Van and Le Thanh, 2022), code-mixing (Bhowmick et al., 2023) and paraphrasing (Marie and Fujita, 2020; Guo et al., 2021a). It is unlikely that back-translation would be an effective method for bolstering the amount of training data for building models for other LRL NLG tasks such as summarisation or dialogue generation. Using back-translation for data synthesis could introduce inconsistencies at a syntactic and semantic level in the training corpus. Given that synthesising code-mixed data has yielded performance increases for translation tasks, it could be insightful to apply this approach to languages that are often interwoven with each other. For example, modern Gaeilge usage could be more accurately modelled if it was trained on a corpus code-mixed to include English, given the strong relationship between both languages (Laoire, 2016). Therefore, it could be interesting to investigate the impact of code-mixing on other NLG tasks. While paraphrasing has produced performance improvements for translation tasks, it has also yielded improvements for other tasks such as question-answering corpus generation and dialogue generation (Sun et al., 2019; Gao et al., 2020; Wu and Wang, 2023). This approach of using paraphrasing models to bolster training data is a promising approach that hinges on the availability of enough data or existing paraphrasing models. More research and experimentation is required to assess the impact of large-scale paraphrasing as a means to augment training data for building NLG models for LRLs. Training language models on a family of related languages has been quite effective at bridging the data gap between LRLs. Given that this language family-based approach can reduce the number of model parameters (Cahyawijaya et al., 2021) and that multilingual modelling can be compared to a 33% increase in training data for an LRL (Chang et al., 2023), it is clear that this is an effective text data augmentation method. It would be interesting to apply this language family-based approach to various groups of related LRLs to further validate this observation.
4.7. RQ7: How Scarce Is the Data in Studies Tackling Low-Resource Language Modelling?
The availability of data for building language models differs considerably between LRLs. Furthermore, it is important to discuss that measuring and comparing the scarcity of data presents a challenge given that the literature often reports different metrics relating to the training data. Some papers clearly document the number of sentences used to train models for LRLs (Agarwal et al., 2022; Mao and Yu, 2024; Pham Van and Le Thanh, 2022; Hong et al., 2024; Gerz et al., 2018). On the other hand, several studies quantify the data in terms of size (Salim et al., 2023; Downey et al., 2024; Acikgoz et al., 2024; Cahyawijaya et al., 2021). Some papers quantify the size of their training dataset by reporting the number of tokens in their respective datasets (Toraman, 2024; Chang et al., 2023; Maimaiti et al., 2021). Moreover, some studies in this review omit metrics relating to the size of their datasets (Niyogi and Bhattacharya, 2024; Holiel et al., 2023). Overall, these varying methods of reporting dataset metrics highlight a limitation in comparing different language modelling approaches for LRLs, given that there is not a universal metric enlisted by these studies that allows us to evaluate and compare each dataset. Although it is a challenge to assess the scarcity across different studies, some insights can be gained from our analysis. Firstly, we propose that there is a sliding scale of data scarcity for LRLs. Specifically, some LRLs such as Khmer were modelled using three different datasets (Agarwal et al., 2022; Mao and Yu, 2024; Pham Van and Le Thanh, 2022), whereas many of languages modelled by the studies in this review were only trained on a single dataset (Berckmann and Hiziroglu, 2020; Li et al., 2024b). While some LRLs such as Uyghur were modelled using a variety of datasets by different studies (Mao and Yu, 2024; Mi et al., 2024), other LRLs have been modelled by a single study using much more data (Gerz et al., 2018). These findings suggest that there is a data disparity within the group of LRLs itself. Secondly, we observe that out of the studies that quantify their data in terms of size, the Turkish dataset is significantly larger than the others (Acikgoz et al., 2024). This reflects the conflicting stance on the resource level of Turkish within the literature in this review (Chang et al., 2023; Toraman, 2024; Yirmibeşoğlu and Güngör, 2023). This contributes to our findings that there is a sliding scale of data scarcity, even among the LRLs themselves.
4.8. RQ8: How Are the Models in These Studies Being Evaluated?
These studies evaluate generative language models for LRLs in many different ways for a broad range of NLG tasks. Similarly to RQ7, the diversity of reported evaluation metrics introduces challenges when comparing results between studies in this review. Given that half of the papers in this review tackle translation, it is unsurprising that BLEU dominates as the most frequently used evaluation metric. However, it is worth highlighting that the BLEU metric suffers from a number of known shortcomings relating to the reporting of results. Post highlights that preprocessing techniques can have a significant impact on BLEU scores, thus impeding on comparability between studies (Post, 2018). Some papers in this review enlist sacreBLEU as a means to address BLEU-based limitations (Wongso et al., 2023; Hong et al., 2024; Liu, 2022). The use of metrics such as ROUGE, chrF, METEOR and relatively newer methods like COMET and sacreBLEU indicates that a wide variety of fine-tuned task-specific models are being trained by studies in this review. The infrequent use of human evaluation methods reveals limitations in most of the studies given that such qualitative insights into model output are important when modelling LRLs. Evaluating whether a model successfully captures the cultural qualities of a language should be essential in determining if an LRL has been modelled accurately. Given that modelling an LRL would be a form of cultural preservation, it would be important to include a qualitative evaluation that ties the research output back to a human context. The lack of a universal approach to evaluating LRL models and the use of flawed and dated, though still prevalent, metrics presents a challenge when comparing the results of different studies within this field of generative language modelling.
4.9. RQ9: What Is the NLG Task?
While a variety of different NLG tasks are modelled by studies in this review, translation dominates as the most prevalent task. Given that this review focuses on modelling LRLs, this dominance could be attributed to the need for translation models as an initial step in digitising and creating models for these languages. Furthermore, the resources for building translation models could be more readily available in contrast with the extensive resources that are required for training other generative models such as dialogue generation models. This observation could further highlight that LRLs are trailing significantly behind more widely spoken languages such as English in terms of the research output for them. Some of the studies contribute a general language model that can be adapted and fine-tuned for different tasks (Gerz et al., 2018; Salim et al., 2023). This type of contribution could be an initial step that could yield further research relating to downstream generative language tasks. Question-answering was also found to be frequently modelled within this review. This research interest could reflect a demand for information retrieval systems for LRLs. Other NLG tasks such as summarisation and dialogue generation were rarely modelled by studies in this review. This could also highlight that LRLs are lacking in research for these downstream tasks. Given the popularity of generative language models, we would expect the frequency of natural language generation tasks to shift from translation to dialogue generation and general language model research in the future.
5. Limitations
This section outlines the considerations and limitations relating to this systematic review. Given that LRLs are a central element of this research, excluding non-English papers could omit relevant literature that addresses the challenge of building generative language models with limited data. Furthermore, the title, abstract and keywords of a paper were examined to identify their relevance during the second pass of the study selection phase. This introduces a limitation in this systematic review as relevant information could have been present in the identified papers but absent from their respective metadata. Although both popular platforms and NLP-specific repositories were searched for relevant studies, relevant studies could exist in less conventional, but still significant sources such as niche linguistic journals or conferences. While the inclusion of synonyms for keywords such as ”low resource language” and ”language model” improves coverage, it does not guarantee that all relevant studies will be identified. In terms of performing the study selection, the use of ASReview as a tool to improve the initial screening process could introduce inaccuracies. The tool does not classify studies but instead predicts their relevance using active learning. Therefore, algorithmic bias could be introduced whereby a reviewer could expect studies seen at the beginning of the selection process to be relevant and could consequently include irrelevant studies. It is worth reporting that a single reviewer worked independently on each stage of the screening process. Although this increases the risk of subjectivity and errors, this reflects the challenges involved with sourcing adequate resources such as multiple reviewers to mitigate bias during the study selection phase. The heterogeneity of results across the different studies imposes a limit on the validity of the comparisons and conclusions between them. Many of the studies report the same metrics such as BLEU, however not all of them include a detailed description of their experimental setup. Moreover, some studies report different or updated forms of metrics (sacreBLEU) for the same generative tasks that make it challenging to compare their results accurately.
6. Conclusion
This systematic review has addressed the challenges posed by data scarcity when building generative language models for LRLs. We have extracted relevant information that contributes to the overall research question by clearly outlining objectives relating to identifying methods for overcoming data scarcity, their impact on model performance and documenting future research directions. By applying a screening process and a series of sub-research questions to each paper identified in the study selection a collection of results was produced.
The results of this systematic review indicate that Indo-European languages dominate LRL modelling research, suggesting that there are disparities within the category of LRLs itself. These findings could suggest that there is a need to define a more nuanced measure of LRL data scarcity. A scale could help to inform a suitable method for overcoming data scarcity. The results also reveal monolingual text data augmentation, multilingual modelling and back-translation were the most commonly used augmentation techniques. Monolingual text data augmentations were found to be effective at improving LRL models, however its efficacy hinges on both the availability and quality of existing language models for neural-based replacement approaches. Therefore, further research is needed to develop NLP models for underrepresented LRLs that currently lack sufficient models, enabling them to benefit from the augmentations discussed earlier. In terms of model architectures, the results indicate that transformer-based models dominate across LRL modelling research. Their popularity can be attributed to their success in modelling a wide range of languages. Although BLEU was the most common evaluation metric, there is a need for universal evaluation metrics that can be used to compare the performance across different LRLs. Universal approaches for reporting the quantity of data used to train LRL models need to be adopted, as this systematic review found that inconsistent reporting made it challenging to compare data scarcity across different studies. Publications in the field of LRL modelling are predominantly driven by ACL and ACM, with a significant number of preprints on arXiv potentially indicating that the field is evolving. The results indicate that translation dominated as the most common NLG task. This could suggest that LRL research prioritises foundational language applications like machine translation.
More research needs to be done to bridge the gap in language inequality and to ensure that less-resourced languages, and communities by extension, are not left behind in the digital age. Developing and releasing conversational language models akin to ChatGPT for LRLs could be a significant step in this direction and consequently requires further efforts. Not only do these LRL models have the potential to preserve entire languages by capturing their linguistic features, but they are also essential for sustaining and promoting language diversity in an age increasingly dominated by large-scale language technologies. The findings of this systematic review aim to support researchers working with low-resource languages by providing a clearer understanding of current strategies for overcoming data scarcity across diverse NLG tasks. Ultimately, this knowledge can contribute to empowering speakers of minority and underrepresented languages through more inclusive technological development.
7. Conflicts of Interest
The authors declare that they have no conflicts of interest related to this work.
8. Acknowledgments
This publication has emanated from research conducted with the financial support of Taighde Éireann - Research Ireland under Grant number 18/CRT/6223.